OpenStudy (anonymous):

use induction to prove the statement counts for all natural numbers n≥1 P(n):1+1/2^2+1/3^2+......+1/n^2≤2-1/n

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

at the start. what i've done is this \[1\div1^2=2+1\div1^2-1\div1\le2-1\div1\] \[1\div t^2=2+1\div t^2-1\div t \ge2-1\div t\] and I'm pretty sure I have done it wrong so far...